Calculating the Direct Materials Price Variance and the Direct Materials Usage Variance

Guillermo's Oil and Lube Company is a service company that offers oil changes and lubrication for automobiles and light trucks. On average, Guillermo has found that a typical oil change takes 25 minutes and 6.2 quarts of oil are used. In June, Guillermo's Oil and Lube had 1,000 oil changes.

Guillermo's Oil and Lube Company provided the following information for the production of oil changes during the month of June:

Actual number of oil changes performed: 1,000
Actual number of quarts of oil used: 6,300 quarts
Actual price paid per quart of oil: $5.40
Standard price per quart of oil: $5.30

Required:

1. Calculate the direct materials price variance (MPV) and the direct materials usage variance (MUV) for June using the formula approach. If required, round your answers to the nearest cent.

MPV $
MUV $

2. Calculate the total direct materials variance for oil for June. If required, round your answer to the nearest cent.

$  

3. What if the actual number of quarts of oil purchased in June had been 6250 quarts, and the materials price variance was calculated at the time of purchase? If required, round your answers to the nearest cent.

What would be the materials price variance (MPV)?

$  

What would be the materials usage variance (MUV)?

$  
Answer #1

Ans:

1) Material price variance = (Standard price-actual price)actual qty = (5.3-5.40)*6300 = 630 U

Material quantity variance = (Standard qty-actual qty)Standard price = (1000*6.2-6300)*5.3 = 530 U

2) Total direct material variance = 630+530 = 1160 U

3)

Material price variance =(Standard price-actual price)actual qty= (5.3-5.4)*6250 = 625 U

Material quantity variance = (Standard qty-actual qty)Standard price (1000*6.2-6300)*5.3 = 530 U